OpenStudy (anonymous):

A cab charges $1.75 for the first mile and $0.25 for each additional mile. Write and solve an inequality to determine how many miles Eddie can travel if he has $15 to spend. Question 6 options: 1) $1.75 + $0.25x ≤ $15; x ≤ 53 miles 2) $1.75 + $0.25x ≥ $15; x ≥ 53 miles 3) $0.25 + $1.75x ≤ $15; x ≤ 8 miles 4) $0.25 + $1.75x ≥ $15; x ≥ 8 miles

OpenStudy (anonymous):

so we know right away its going to be 0.25x + 1.75

OpenStudy (anonymous):

if he has $15 to spend 1.75+0.25x = 15 15 - 1.75 = 13.25

OpenStudy (anonymous):

0.25x = 13.25 13.25/0.25 = 53

OpenStudy (anonymous):

X cannot be greater then 53

OpenStudy (anonymous):

if x was for example 54, wed get 1.75+54(0.25) = 15.25, and he only has $15 to spend